Consider a pair of ten sided dice. (a) what is the probability of rolling nines for both dice. (b) what is the probability that the sum of the dice is 18?
1
Expert's answer
2020-09-22T16:02:14-0400

Solution to a :

Probability of rolling 9 and 9:


"P(x=9 and x=9)= p(x=9) * P(x=9)"

And

"P(x=9) = {1 \over 10}"

Therefore ;

"= ({1 \over 10})({1 \over 10}) = { 1 \over 100}"

Answer: 0.01


Solution to b

Probability sum is 18:


"=P(9 and 9) or P(10 and 8) or P(8 and 10)"

"= P(x=9) * P(x=9) +"


"P(x=10) * P(x=8) + P(x=8)*P (x=10)"


"= ( {1 \over 10})( {1 \over 10}) + ({1 \over 10}) ({1 \over 10}) + ( {1 \over 10})( {1 \over 10})"

"={1 \over 100} + {1 \over 100} + {1 \over 100} = {3 \over 100}"

Answer: 0.03